OpenStudy (anonymous):

Solve for x using the natural logarithm:

OpenStudy (jdoe0001):

kinda hard to do, when the expression is invisible

OpenStudy (anonymous):

\[2\times4^{x} = 9.5e ^{-2x}\]

OpenStudy (anonymous):

Thank you for your patience, @jdoe0001

OpenStudy (anonymous):

@SithsAndGiggles Can you help?

OpenStudy (jdoe0001):

hmmm can you pot a quick screenshot of the material?

OpenStudy (anonymous):

@zepdrix Can you help? I'm desperate :(

zepdrix (zepdrix):

:D

zepdrix (zepdrix):

Oh you left :c

OpenStudy (anonymous):

I'm back!

zepdrix (zepdrix):

:D

OpenStudy (anonymous):

:D

zepdrix (zepdrix):

\[\large\rm 2\cdot4^{x} = 9.5\cdot e ^{-2x}\]I don't like decimals, fractions are better, so I'm gonna do something a little sneaky. Let's umm.... let's multiply both sides by 2, then divide each side by 4,\[\large\rm 4^x=\frac{19}{4}e^{-2x}\]

zepdrix (zepdrix):

Then take natural log each side I guess, ya? :)

zepdrix (zepdrix):

\[\large\rm \ln\left(4^x\right)=\ln\left(\frac{19}{4}e^{-2x}\right)\]We have to apply a bunch of fun log rules to solve for x.

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\log(a^b)=b\cdot \log(a)}\]Do you see how this orange rule might help us on the left side of our equation?

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

Give me one sec

OpenStudy (anonymous):

Wait why doesn't 4 cancel out when you divide it?

zepdrix (zepdrix):

Umm if that was too confusing, we can just do it this way instead :)\[\large\rm 2\cdot4^{x} = 9.5\cdot e ^{-2x}\]Dividing by 2 gives us:\[\large\rm 4^{x} = 4.75\cdot e ^{-2x}\]I just didn't like the decimal value, so I changed it to a fraction, but we can leave it like this if it's easier to understand.

OpenStudy (anonymous):

Ok

zepdrix (zepdrix):

So taking log of each side gives us\[\large\rm \ln\left(4^x\right)=\ln\left(4.75\cdot e^{-2x}\right)\]

OpenStudy (anonymous):

So you get x(ln4) = ln(4.75e^-2x) ?

OpenStudy (anonymous):

And now ln and e cancel each other out

zepdrix (zepdrix):

Left side looks good. Woops! Don't try to cancel out the e just yet! Gotta do some more simplifying before you can do that.

zepdrix (zepdrix):

\[\large\rm \color{orangered}{x\ln4}=\ln\left(4.75\cdot e^{-2x}\right)\]How bout the right side? How can we apply this blue rule?\[\large\rm \color{royalblue}{\log(a\cdot b)=\log(a)+\log(b)}\]

OpenStudy (anonymous):

x(ln4)= ln(4.75) + ln(e^-2x) ?

zepdrix (zepdrix):

Ok great! Now we don't have the number in front of the e, so we can "cancel" them out like you wanted.

OpenStudy (anonymous):

xln4= ln4.75 - 2x ?

zepdrix (zepdrix):

\[\large\rm x \ln4=\ln4.75+\ln e^{-2x}\]\[\large\rm x \ln4=\ln4.75-2x\]mmmm k good!

zepdrix (zepdrix):

Let's get all of our x's to one side and try some factoring.

OpenStudy (anonymous):

x(ln4) + 2x= ln4.75 ?

zepdrix (zepdrix):

cool :) the terms on the left side both have something in common. try to factor it out :O

OpenStudy (anonymous):

x(ln4 + 2)= ln4.75 ?

zepdrix (zepdrix):

Good good good, how you gonna wrap it up? :)

OpenStudy (anonymous):

x = 0.4601 ?

zepdrix (zepdrix):

Yayyy good job \c:/ If your teacher wants you to leave it as an exact value you would have:\[\large\rm x=\frac{\ln4.75}{2+\ln4}\] But yes, that's a correct decimal approximation.

OpenStudy (anonymous):

Yay!!! You are the BEST!!! Thanks so much!